OpenStudy (anonymous):

evaluate the following limit \[\lim_{x \rightarrow 0^{+}}\cot x-\frac{ 1 }{ x }\]

OpenStudy (anonymous):

I know that it is infinity-infinity but I don't know how to change it into a problem I can work with

terenzreignz (terenzreignz):

Well... How about this... \[\huge \cot \ x - \frac1{x}=\frac{x\cos \ x-\sin \ x}{x\sin \ x}\]Now it's just a fraction. L'hopital away, champ ^.^

OpenStudy (anonymous):

wow, thanks

terenzreignz (terenzreignz):

o.O Here I am again, overcomplicating things... maybe you could try this, it might be a lot less headache-inducing... \[\huge \cot x-\frac1{x}=\frac{x\cot \ x -1}{x}\]

terenzreignz (terenzreignz):

Sorry about the first one, this one only requires one use of L'Hopital...

terenzreignz (terenzreignz):

Or then again, maybe not. I don't know anymore O.o

OpenStudy (anonymous):

well when I did the first one i simplified it down to -sinx/cosx and substituted zero and got 0 which is the correct answer in the book. But when you're second guessing yourself I want to make sure I did it correctly

terenzreignz (terenzreignz):

Never mind the second... I tried... it loses more than it gains... it gets a cot x in the numerator, which is troublesome. Stick to my first idea :) Instincts are to be trusted, after all :)

OpenStudy (anonymous):

lol alright i get it, happens to me all the time.

terenzreignz (terenzreignz):

Good hunting :)